Subj: HST REPORT #886 Note: Both the AP-5 and AP-7 computers where down with problems on the operational system disk. This made the ALLIN1 and MAIL functions inoperative on Wednesday May 12, 1993, and Thursday May 13 causing a delay in the distribution of the daily report via the electronic mail. HUBBLE SPACE TELESCOPE DAILY REPORT #886 PERIOD COVERED: 0800 EST Daily Status Report as of 133/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FOS/BL 4697 (FOS Polarimetric Calibration) The Faint Object Spectrograph was used to make a total of 17 calibration observations on both the red and blue sides to obtain measurements of the internal Pt-Ne lamps. All seventeen observations were received and displayed. They were all reported to look okay. No unexpected problems were noticed. 1.2 Completed GHRS/2 3945 (G160M Spectra of Eta Car S Condensation - Cycle 2) The High resolution Spectrograph was used to make spectral observations of a gaseous knot in the southern part of the nebula Eta Carnae. In the first set of eleven observations, all were received and displayed. They were reported to look okay. There were two carrousel resets events which occurred, and generated status buffer error messages. The second carrousel reset delayed the final wavelength calibration until one minute after the scheduled start time of the final science observations. This observation did not occur. 1.3 Completed Two Sets of FOC/96 3801 (Search for QSOs suitable for Subsequent Observations of He II) The Faint object Camera made ultraviolet images of five different quasars (2139-4434, 2050-359, 2049-353, 2038-371, and 1601+3754) to observe the level of their UV flux. Two recentering events occurred during two separate observations, and no other anomalies were noted. 1.4 Completed WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Wide Field Camera was used to make two dark calibrations images. These were pre-flashed exposures which lasted 1200 seconds each. Both images were received and displayed. No anomalies were noted. 1.5 Completed FOS/BL 4054 (Probing the Nuclear Regions of Seyfert Galaxy NGC-5548) The Wide Field Camera was used to make observations of Seyfert galaxy NGC-5548 to look for time variations in its continuum, and emissions in Helium II and Carbon IV. The observations were executed as scheduled, and no problems were reported. 1.6 Completed WFC 4381 (HST Medium-Deep Survey Cycle 3) The Wide Field Camera was used to make two high latitude medium-deep sky survey observations taken in parallel with the prime science observations. The OSS monitored the observations, and assessed the data quality. No problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 4 The acquisition at 132/1443Z failed to attain fine lock, but did obtain course track. Three subsequent reacquisitions also failed fine lock, but obtained course track. Scheduled Reacquisitions : 4 Successful : 4 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Significant HSTARs: None 2.4 Operations Note: The GSFC Internal Simulation (Type 1 + Replan) was conducted on Wednesday, May 12, and Thursday May 13. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FOS/RD 4817 (Quasar Absorption Line Survey: Cycle 2 Observations) 3.2 Complete FOC/96 3504 (High Spatial Resolution Imaging and Spectroscopy of AGNs: Cycle 2) 3.3 Complete WFC 4849 (GTO High Latitude Parallel Program: Cycle 3 High) 3.4 Complete FOC/96 3801 (Search for QSOs Suitable for Subsequent Observation of He II) 3.5 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.6 Complete GHRS/2 3664 (Probing the Gas of the Superbubble LMC2) 3.7 Complete FOS/BL 4054 (Probing the Nuclear Regions of Seyfert galaxy NGC-5548) 3.8 Complete WFC 4381 (HST Medium-Deep Survey - Cycle 3) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-